A random sample found that thirty-six percent of 50 Americans were satisfied with the gun control laws in 2017. Compute a 97% confidence interval for the true proportion of Americans who were satisifed with the gun control laws in 2017. Fill in the blanks appropriately. A 97% for the true proportion of Americans who were satisfied with the gun control laws in 2017 is(_____)(_____)(Keep 3 decimal places)
Answer #1

\alpha = 1 - 0.97 = 0.03

From Z table,

Z\alpha/2 = 2.1701 (For two tailed confidence interval)

97% confidence interval for p is

[ \hat{p} - Z\alpha/2 * sqrt ( \hat{p} ( 1 - \hat{p} ) / n ] < p < [ \hat{p} + Z\alpha/2 * sqrt ( \hat{p} ( 1 - \hat{p} ) / n ]

0.36 - 2.1701 * sqrt [ 0.36 * ( 1 - 0.36) / 50] < p < 0.36 + 2.1701 * sqrt [ 0.36 * ( 1 - 0.36) / 50]

0.213 < p < 0.507

97% CI is ( 0.213 , 0.507 )